OpenStudy (anonymous):

HELP ASAP! (: Thanks 31. Find the tenth term of the pattern. 0, 1, 4, 9, 16, 25, 36, . . . A. 64 B. 81 C. 100 D. 49

OpenStudy (kirbykirby):

They are all perfect squares... \(0^2=1 \\1^2 = 1\\2^2=4\\ 3^2=9\\ 4^2=16\\ 5^2=25\\...\)

OpenStudy (anonymous):

so 10*10=100 C?

OpenStudy (anonymous):

the answer is b due to 0, 1, 4, 9, 16, 25, 36, 49 , 64, 81 1 3 5 7 9 11 13 15 17

OpenStudy (kirbykirby):

Yes it would be B, just be careful that the 10th term is actually \(9^2=81\) (if you count the 0 term, there are ten terms from 0 to 9, :)! )

OpenStudy (anonymous):

oh ok